A random sample of 10 resistors is to be tested. From experience, it is known that the probability of a given resistor being defective is 0.057. Let X be the number of defective resistors. Requirements: a) What kind of distribution function would be recommended for modeling the random variable X? b) According of distribution function in (a), what is the probability that in the sample of 10 resistors, there are more than 1 defective resistors in the sample?
Answer by Theo(13342)

i think possibly a binomial probability distribution type of problem.
n = 10
x = 0 to 10
p = .057
q = 1-p = .943

p(x) = p^x * q^(n-x) * c(n,x)

p(x > 1) = 1 minus p(0) minus p(1).

p(0) = .057^0 * .943^10 * c(10,0) = .5560539464
p(1) = .057^1 * .943^9 * c(10,1) = .3361089602

p(0) + p(1) = .891629066

1 - (p(0) + p(1) = .1078370934
that should be your solution, assuming this is a binomial probability distribution type of problem.

here's the complete set of probabilities for x from 0 to 1.
the sum is 1 as it should be.